Tech Staff Meeting Notes January 7, 2008

Current Top Priorities are: Station maintenance and GPS wrapup.

There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about

The task list was updated October 9, 2007 to incorporate some tasks previously performed by Greg. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything.


Station Maintenance, NMSZ

Station maintenance is the top priority.

Holly's list on 1/7/2008: TIPT


HICK.HHN - drifting

Vandalism update. MORT is rebuilt and back. GRAT is still down. Need to look for a spot on NWCC property to move TIPT.

Need a dual mux assembly or two. Have the AMUX64T's. Pat will build up the rest of the hardware.
MKTA spare air conditioner is repaired. Needs to be returned to the doghouse.
Need a remote reset at lnxt. We'll use an ADAM 6060 internet reset device.
Need new RF preamps, radio align, etc, etc at MKTA.
PEBM and HENM sometimes have telemetry fades. Should reevaluate comms.
Brush at NMAD isn't too bad, we may just need to trim up around the guy anchors.

Mud Island vertical array is operational.

The recorder disks should be reformatted as a routine maintenance task next time access is available. The constant read/write to the fat file systems lead to fragmentation. Don't ya just love dos?
Lost contact with recorders 1 and 3. Can reset all 7 recorders which implies the equipment enclosure at each node has power and that at least the reset radio and telemetry are working.


We're in the process of getting James out of the station maintenance business. GPS stations will be treated as any of the seismic network stations. Mac will be the lead for routine maintenance. Data handling and processing is still done by Bob et al.

Greg is in charge of getting anything done to the GPS stations to Steinerize them. This is necessary to make sure we can maintain them without burning alot of time on reapplying bandaids. Note that they are in pretty good shape already. The primary hassle is communications.

Most of the work for the upgrades and two new stations is complete. The telemetry link between NMKM and NMAD should work but doesn't. Jim and Steve may have found the problem. Left to do is the communication and the cages.

This graphic shows the telemetry and data flow. The dashed lines are plan B in the unlikely event that the problems with the NMAD-NMKM link can't be resolved.

Items left:
Comms for MIAR, CJTR.
Install New Madrid end of link (4' dish is hung, needs aligned).
RLAP needs solar regs.
The moxa fiber to cat5 converter is unstable at RLAP. Can't reproduce in shop.
Camp Robinson modem is down, possible lighting.
Portageville needs a visit to check router enclosure for fan and general SOH. Also need to replace 6A power supply with a 3A.


Station Maintenance

Murphy repair seems to be on again off again. Mitch will ask Alena and Mark about it.
BRBC has a new XLAK GPS antenna. Mitch adjusted the minimum time for an alarm. Still get them. Not sure which of the dozen possible sources is actually generating the alarm. A handheld gets plenty of satellites next to the building. A multipath problem at that particular spot? Could also just be the clock settings: all the others had to be adjusted, did this one escape?
Chestnut Mt interference magically disappeared. I hate that.
Star Mt remote reset isn't working properly. We currently only have power cycling capabilities for the worm machines at both Athens and Starr Mt. The problem in both cases appears to be with the reset card/hardware at the Athens end.
John can't log into smeagol. He'll get with Steve or Al to figure out why.
CVTN and SWET have leaky freewave enclosures. This is a common problem with that style enclosure when mounted sideways.

Network Completion

The map shows stations with yellow triangles being new ones.
SMNC4done (see note above)
BHT8move (eventually)

Routine Maintenance Kudzu, battery check, etc, etc.

Need to swap out the old node computers at NHIN, GNAR, and AZTN. Steve needs to go to NHIN (but Steve isn't allowed to do anything else until the ew 7.2 upgrade is complete).

Second Tier Priorities

CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready.

Steve spoke with Dean Kemp at NWS. They are willing to let us piggyback on their cable internet (not Dean is looking into the possibility of a static IP. This would get NAIT online. Bill will call to check on this.

Lenox Agreement. Mitch spoke with Ron Martens, EMF Broadcasting. There is a desire for some sort of MOU and a visit to CERI. Mitch will set this up.

FDL. This is the FDL Policy. .
Numerous cables from the vandalism need repaired.
Need 2 firewire cables. They are $500 each from the manufacturer and generic ones don't appear to work. The speed isn't sufficient. So what's special about the expensive ones? Bill will try to find out.

As Time Permits

3876 Power The power for the new lighting on the east side of the general parking lot to the west of CERI is coming from 3876. Chris will check it out to assess the added lightning hazard and see if anything can be done about it. Might be hard to do without the wiring diagram. Chris will make one if needed as time permits.

Have a water fountain for 3876. Needs to be installed.

3890 Power Chris is going to make a diagram for 3890. Good luck son.

Inventory Jim has all the broadband sensors defined and is working on additional equipment (e.g. isis) and the connections. Mitch will install Bill's account in mysql and inv so he can start building stations. Mitch needs to sit down with Jim and figure out what we need to have ISTI help us with. Sid will be available after about mid-May to finish up anything we need his help to take care of. Apparently the downloaded docs from guralp for the cmg5td's have the old +-4g contants. Mitch has paper documentation with the correct +-2g constants.

Steiner Brain Dump Seminar Need more of them as weather permits. Especially calibration and computer room power, general station power, and shake table operation.


Need new doors on the USGS shed. A small child with a plastic screwdrive could get in.

Withers needs to determine current orientation for MPHB so we can pull it and orient correctly.

Need to teach someone (our hazmat dude?) to neutralize etchant and clean the etching tanks. Along those same lines, Greg has a ton of drawings that play only with the "Eagle" software on his computer. Mitch is unaware of a way to convert to a more generic format. Also require a printer that does negative (the one on Greg's serial port does, others do not). Mitch knew this dependence on vendor specific software would bite us one day. That day is soon.

Pink Palace hardware needs to be rebuilt.

John needs a spectrum analyzer.

Evaluate computer room power. Steve will be the keeper of the spreadsheet post-Steiner.

Build more battery charge controllers.